Mei earn $2 less than three times as much as Olivia. Shane earns twice as much as mei. Together they earn $594 per week. How much does each person earn per week?
How to solve ?

Answer by lwsshak3(11628)

Mei earn $2 less than three times as much as Olivia. Shane earns twice as much as mei. Together they earn $594 per week. How much does each person earn per week?
***
let x=amt Olivia earns
3x-2=amt Mei earns ($2 less than three times as much as Olivia).
2(3x-2)=amt Shane earns (twice as much as mei)
..
x+3x-2+2(3x-2)=594
x+3x-2+6x-4=594
10x=600
x=60
3x-2=178
2(3x-2)=356
amt Olivia earns=$60
amt Mei earns=$178
amt Shane earns=$356
..